Use the following information about the overhead reach distances of adult females: ? = 205.5cm, ?=8.6 cm, and overhead reach distances are normally distributed.

A random sample of 50 adult females is selected. Find the probability that the mean overhead reach of the sample is between 185 cm and 220 cm.
Answer #1

Ans:

Given that

mean=205.5

standard deviation=8.6

n=50

z=(185-205.5)/(8.6/sqrt(50))=-16.86

z=(220-205.5)/(8.6/sqrt(50))=11.92

P(-16.86<z<11.92)=P(z<11.92)-P(z<-16.86)=1-0=1.0000